### 05 Dec 2023 v2

# Robust Parameter Estimation for Rational Ordinary Differential Equations

Abstract
We present a new approach for estimating parameters in rational ODE models from given (measured) time series data.
In a typical existing approach, one first tries to make a good initial guess for the parameter values. Then, in a loop, the corresponding outputs are computed by solving the ODE numerically, followed by computing the error from the given time series data. If the error is small, the loop terminates and the parameter values are returned. Otherwise, heuristics/theories are used to possibly improve the guess and continue the loop. A downside of this approach is non-robustness, as there are no guarantees for the result of the loop iterations to be predictably close to the true parameter values.
In this paper, we propose a new approach, which does not suffer from the above non-robustness. In particular, it does not require making good initial guesses for the parameter values. Instead, it uses differential algebra, interpolation of the data using rational functions, and multivariate polynomial system solving, and has a potential for a complete user control over the error of the estimation (the actual error analysis is left for the future research). We also compare the performance of the resulting software with several other estimation software packages.
Keywords Parametric ODE Models  $\cdot$ Parameter Estimation  $\cdot$ Differential Algebra  $\cdot$ Symbolic-Numeric Differentiation  $\cdot$ Mathematical Software
1 Introduction
1.1 Overall problem
Parametric ODEs are ubiquitous in science and engineering. In order to use them, one first needs to estimate the parameters, which is usually done using measured time series data. Hence, the overall problem is: given a parametric ODE and some measured data, to estimate the parameters.
1.2 State of the art
Due to its importance, this problem has been the subject of intensive research efforts that have yielded various theories and methods. They can be roughly categorized into three approaches: shooting, two-stage, and algebraic.
Shooting approaches (also known as simulation-based approaches) directly use the ODE solver to compute the loss function of the current estimate. This is used to define a nonlinear optimization to improve the parameter estimates. This popular approach has seen many software implementations; to list a few: AMIGO2, COPASI, Data2Dynamics, SBtoolbox2/IQM; see [3, 5, 38, 65, 58, 56, 36, 29, 23, 21, 2, 54, 25, 63, 19]. Many studies have further developed the shooting methods to improve robustness, adding techniques such as multiple shooting [10, 1, 60], automatic differentiation and adjoint techniques for more accurate derivatives [41, 46, 33], and landscape smoothing techniques like the prediction-error method [34], all of which help the optimization techniques avoid local minima.
Two-stage approaches (also known as collocation, smoothing, or principle differential analysis) avoid numerically solving the ODE by fitting collocation polynomials to the data to obtain predictions of the data which are then directly fit against the data [52, 14]. As the result, these methods are generally much more efficient. Previous studies have used various mathematical objects such as B-splines [13, 48, 35] and local least-squares [37, 55, 53, 15]. One downside to two-stage methods is requiring observation of all states. Furthermore, it has been noted that the error introduced through the derivative estimation tends to lead to less accurate final results in real-world scenarios with less dense data [48, 14]. These studies suggest finishing a parameter estimation by using the output of a two-stage method as the initial condition to a simulation-based method.
Algebraic approaches tackle rational differential models by exploiting theoretical results from differential algebra. The paper [23] expresses parameters algebraically in terms of derivatives of input and output functions and then estimates these derivatives using the numerical data. Two difficulties with this approach are the size of these algebraic expressions and the computational complexity in obtaining them. The papers [11, 31, 62, 67] study how to use input-output equations in parameter estimation and how to lower the orders of derivatives used in the estimation (because higher order derivatives typically introduce larger errors with numerical data) by different approaches, including using integrals, cf. [57, Section 3.2]. Computing input-output equations could be computationally very costly, though recently there have been improvements in algorithms computing them [20]. The paper [49] considers an algebraic approach to parameter estimation in linear systems. Finally, the paper  [32] investigates the parameter manifold in the case where the differential model is structurally non-identifiable.
1.3 Our approach and novelty
Our approach could roughly be classified as a two stage approach. However, it is substantially different from the previous two stage approaches, in that it computes optimal parameters robustly. For instance, it can reliably handle even degenerate cases such as when the parameters are only locally identifiable. It achieves the robustness by judiciously combining:
  • differential algebra approach from the parameter identifiability analysis developed in [27, 28],
  • estimation of the derivatives of the ODE solutions from rational interpolation of the data,
  • overdetermined polynomial system solving by squaring the system and filtering the solution set.
Our approach has the following novel features:
  1. 1.
    Our approach does not involve making initial guesses or global optimization; it instead relies on differential algebraic properties of the ODE model to propose candidates for parameter values.
  2. 2.
    Out of the possible solution candidates, the procedure computes the best fit for the given data including cases in which such fit may not be unique. That is, accounting for possible local identifiability of the model, we report all valid parameter values. The user can then quickly filter out some (or all) of the not possible parameter values if the users has additional information about these values such as ranges they take values in.
As such, this technique can be used in isolation for dense data, or together with shooting methods by providing more robust initial conditions for sparse data.
1.4 Software Implementation
We implemented our approach into a Julia package called “ParameterEstimation.jl”. All code and benchmarking data are available at https://github.com/iliailmer/ParameterEstimation.jl.
2 Problem Statement
In this section, we precisely state the problem solved by our approach and illustrate it using a toy example. We also show how to enter/solve the problem using our soft implementation.
Input:
  1. 1.
    An ODE model $\Sigma$
    $\begin{cases}\bm{x}^{\prime}(t)=\bm{f}(\bm{x}(t),\bm{u}(t),\bm{\mu})\\ \bm{y}(t)=\bm{g}(\bm{x}(t),\bm{u}(t),\bm{\mu})\\ \bm{x}(0)=\bm{x_{0}}\end{cases}$ (1)
    where we use bold fonts for vector:
    • $\bm{f}$ and $\bm{g}$ are rational functions describing the model,
    • $\bm{x}$ are state variables,
    • $\bm{u}$ are input (control) variables,
    • $\bm{y}$ are output variables,
    • $\bm{\mu}$ and $\bm{x_{0}}$ are unknown parameters.
  2. 2.
    A data $D=((t_{1},\bm{y}_{1}),\ldots,(t_{n},\bm{y}_{n}))$ where $\bm{y}_{i}$ is the measured value of $\bm{y}$ at time $t_{i}$
Output: Estimated values for the parameters $\bm{\mu}$ and $\bm{x_{0}}$.
Input: (Toy example)
$\begin{array}[c]{lll}\Sigma&:&\left\{\begin{array}[c]{lll}x^{\prime}&=&-\mu x\\ y&=&x^{2}+x\\ x\left(0\right)&=&x_{0}\end{array}\right.\\ D&:&\left(0.000,2.000\right),\left(0.333,1.563\right),\left(0.666,1.229\right),\left(1.000,0.974\right)\end{array}$
Output: $(\mu,x_{0})\;\approx\;(0.499,1.000)$
The result above is good because we have “simulated” the measured output values $D$ by numerically solving the model $\Sigma$ for $\mu=0.500$ and $x_{0}=1.000$.
To solve the toy example problem using our software, enter the following lines into Julia console:
using ParameterEstimation
using ModelingToolkit
# --- Model
@parameters mu
@variables  t x(t) y(t)
D = Differential(t)
@named Sigma = ODESystem([D(x) ~ -mu * x],t,[x],[mu])
outs = [y ~ x^2 + x]
# --- Data
data = Dict(
  "t"     => [0.000, 0.333, 0.666, 1.000],
  x^2 + x => [2.000, 1.563, 1.229, 0.974])
# --- Run
res = estimate(Sigma, outs, data);
You should see the following output:
Parameter(s)        : mu   = 0.499
Initial Condition(s): x(t) = 1.000
3 Our Approach
We will briefly describe our approach by first giving a flow chart of the steps.
Input: $\begin{array}[]{ll}\text{Model:}&\left\{\begin{array}[]{ll}x^{\prime}&=-\mu x\\ y&=x^{2}+x\end{array}\right.\\ \text{Data:}&\begin{array}[]{l}\left(0.000,2.000\right),\left(0.333,1.563\right)\\ \left(0.666,1.229\right),\left(1.000,0.974\right)\end{array}\end{array}$ of values for the parameters = 1 Differentiated system: $y$ $=x^{2}+x$ $x^{\prime}$ $=-\mu x$ $y^{\prime}$ $=2xx^{\prime}+x^{\prime}$ $x^{\prime\prime}$ $=-\mu x^{\prime}$ $y^{\prime\prime}$ $=2\left(x^{\prime}x^{\prime}+xx^{\prime\prime}\right)+x^{\prime\prime}$ Polynomial system: 2.002.00 2.002.00 $\approx x_{0}^{2}+x_{0}$ $x_{1}$ $=-\mu x_{0}$ $-1.50$ $\approx 2x_{1}x_{0}+x_{1}$ $x_{2}$ $=-\mu x_{1}$ 1.221.22 1.221.22 $\approx 2\left(x_{1}^{2}+x_{0}x_{2}\right)+x_{2}$ Solutions: $S_{1}=\left\{\begin{array}[]{ll}\mu&=0.499\\ x_{0}&=1.000\end{array}\right.S_{2}=\left\{\begin{array}[]{ll}\mu&=0.249\\ x_{0}&=-2.000\end{array}\right.$ Errors: $\begin{array}[]{ll}S_{1}\Rightarrow{\bf 6.87\cdot 10^{-4}}&S_{2}\Rightarrow 2.22\cdot 10^{-2}\end{array}$ Output: $\mu=0.499,~{}x(0)=1.000$ 1) Find # values for parameters 2) Differentiate 3) Use the data 4) Find all solutions5) Compute errors6) Select
Now we elaborate on each step in the above flow chart.
  1. 1.
    Use the model $\Sigma$ to find the number of values for parameters (see [27, 28] for details, cf. [4]).
    Toy example: There is only one value for the parameters.
  2. 2.
    Differentiate the model $\Sigma$ sufficiently many times (see [27, 28] for details).
    Toy example:
    $\left\{\begin{array}[c]{lll}y&=&x^{2}+x\\ y^{\prime}&=&2xx^{\prime}+x^{\prime}\\ y^{\prime\prime}&=&2\left(x^{\prime}x^{\prime}+xx^{\prime\prime}\right)+x^{\prime\prime}\\ x^{\prime}&=&-\mu x\\ x^{\prime\prime}&=&-\mu x^{\prime}\end{array}\right.$
  3. 3.
    Use the data $D$ to derive a numerical overdetermined polynomial system from the above equations.
    1. (a)
      Approximate the output functions $y$ by interpolating the data $D$ into rational functions $g$.
      Toy example:
      $\begin{array}[]{ll}y&\approx\;g\;=\;\frac{0.58t^{2}-3.11t+6.82}{t+3.41}\\ \end{array}$
    2. (b)
      Approximate the system, obtained from Step 2), by replacing $y$ with $g$.
      Toy example:
      $\left\{\begin{array}[c]{rll}\frac{0.58t^{2}-3.11t+6.82}{t+3.41}&\approx&x^{2}+x\\ \frac{0.58t^{2}+3.96t-17.41}{\left(t+3.41\right)^{2}}&\approx&2xx^{\prime}+x^{\prime}\\ \frac{48.32}{\left(t+3.41\right)^{3}}&\approx&2\left(x^{\prime}x^{\prime}+xx^{\prime\prime}\right)+x^{\prime\prime}\\ x^{\prime}&=&-\mu x\\ x^{\prime\prime}&=&-\mu x^{\prime}\end{array}\right.$
    3. (c)
      Evaluate the above system at $t=0$.
      Toy example:
      $\left\{\begin{array}[c]{rll}2.00&\approx&x_{0}^{2}+x_{0},\\ -1.50&\approx&2x_{1}x_{0}+x_{1},\\ 1.22&\approx&2\left(x_{1}x_{1}+x_{0}x_{2}\right)+x_{2}\\ x_{1}&=&-\mu x_{0}\\ x_{2}&=&-\mu x_{1}\end{array}\right.$
      where a shorthand notation $x_{i}=x^{\left(i\right)}\left(0\right)$ is used. Note that the system consists of $5$ equations in 444~{}4unknowns: $\mu,x_{0},x_{1},x_{2}$.
  4. 4.
    Find all the solutions of the over-determined system of polynomial equations. For this, first we make the system square. Then we find all the real solutions of the square system (for instance, using homotopy continuation method, see [12]).
    Toy example:
[…]
  5. 5.
    Compute errors of the solutions.
    For each solution $\left(\widehat{\mu},\widehat{x}_{0}\right)$
    1. (a)
      Compute $\widehat{y}\left(t_{i}\right)$ by solving $\Sigma\left(\widehat{\mu},\widehat{x}_{0}\right)$ using numerical ODE solver.
    2. (b)
      Compute error $e$ between the computed values of $\widehat{y}$ and the measured values of $y.$
    Toy example:
[…]
  6. 6.
    Select $k$ solutions with smallest error, where $k$ is the number of parameter values from Step 1). If the user has a priori information about parameter ranges, we can first filter the solutions before choosing the $k$ best solutions.
    Toy example: $(\mu,x_{0})\;\approx\;(0.499,1.0000)$
4 Implementation details
Steps 1) and 2)
The two steps are are done by calling SIAN [27, 28].
Step 3)
We perform interpolation of the dataset $D$. In this interpolation, we have considerable freedom in choosing the interpolation method, with the following desired properties:
  1. 1.
    easily calculated analytic derivatives, which are as well interpolated as the function itself
  2. 2.
    the ability to fit a wide variety of shapes, including asymptotic behavior, decay to 0, and periodicity,
  3. 3.
    a preference for parsimonious models, and
  4. 4.
    numerical stability.
Our software supports a variety of modern interpolation methods, via BaryRational.jl and other interpolation packages. Of note, we support Fourier extension, Floater-Hormann rational interpolation [24], as well as the “AAA" algorithm introduced in [45], which uses a barycentric rational representation adapted to input data. Experimental data show that AAA rational interpolation interpolates smooth functions in such a way that, for a wide variety of data outputs from differential equations, even higher order derivatives are well interpolated. In our program, we use AAA as a default interpolation scheme, but allow iterating over multiple schemes in a parallel fashion to select the best one.
Once interpolation is done, we apply high-order automatic differentiation provided by the TaylorSeries.jl [7] and ForwardDiff.jl packages to the interpolated function to estimate derivative values of higher order.
We substitute the estimated values of the derivatives of the output $y$ into the system obtained in Step 2). This typically results in an over-determined system of polynomial equations.
Step 4)
Since the obtained polynomial system is typically overdetermined, the next step is to square the system carefully. This is achieved by collecting equations one by one, given that a newly added equation increases the rank of the Jacobian of the system. If, for a given equation, the rank is unchanged, the equation is not added. Then we find all solutions of the squared system. By default, we are using homotopy-based solving from HomotopyContinuation.jl. We also support MSolve [9] via Oscar algebra system in Julia [47, 18]. This method finds all solutions with theoretical guarantees. However, it has been less stable for some of the systems we encountered. All results presented here use homotopy-based solving, the default option.
Steps 5) and  6)
For each obtained solution (parameter estimates), we compute the data by numerically solving the model. Then we determine errors between the computed data and the given (measured) data. We keep the estimates with minimal errors. Such estimates may exist for different interpolation schemes; we keep the ones with minimal error across all schemes.
Extra features
We would like to emphasize the following extra features:
  • The iteration over interpolation schemes and underlying polynomial solving via HomotopyContinuation.jl [12] are parallelizable and can take advantage of a multicore (or multithreaded) computing environment.
  • ParameterEstimation.jl provides identifiability interface to SIAN [27, 28]. One can assess local and global identifiability of the parameters and initial conditions before estimation.
Current Limitations
We list a few limitations of the current implementation:
  • As a two-stage method, the issues common to this class of methods apply. The data must be time series data which is sufficiently dense for the rational polynomial fitting to give accurate derivative approximations.
  • Our program supports rational ODEs models (the right-hand sides are rational functions), while software such as AMIGO2 and IQM also support ODE models that are not necessarily rational.
  • We use SIAN that is based on a Monte-Carlo algorithm that returns the bound on number of solutions with high probability, though detailed analysis is still ongoing.
  • The high order derivative approximations via rational polynomials may suffer from errors due to noise or rounding. We leave the extension of the technique to usage with robust derivative approximations [35, 16] as a topic of further study.
5 Performance
Median of Relative Errors in %
Software IQM SciML AMIGO2 Parameter
Estimation.jl
Search Range [0,1] [0,2] [0,3] [0,1] [0,2] [0,3] [0,1] [0,2] [0,3] Any
Harmonic, eq. 2 64.2 63.0 65.1 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Van der Pol, eq. 4 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
FitzHugh-Nagumo, eq. 6 68.8 139.2 159.7 0.6 0.2 0.3 0.0 0.0 0.0 0.0
HIV, eq. 8 75.5 89.7 133.8 2.0 7.9 57.8 0.0 0.0 0.0 0.0
Mammillary 3, eq. 10 74.3 88.1 85.6 5.6 12.8 11.8 0.0 0.0 0.0 0.0
Lotka-Volterra, eq. 12 71.7 76.5 76.5 22.1 55.0 75.6 0.0 0.0 1.0 0.0
Crauste, eq. 14 85.3 109.3 139.8 26.5 67.2 144.8 0.0 77.6 24.5 0.0
12
Biohydrogenation, eq. 16 81.2 113.4 78.9 53.9 110.6 284.6 0.0 0.0 22.2 0.0
Mammillary 4, eq. 18 76.0 88.4 105.1 46.9 61.1 97.1 0.0 53.8 51.9 0.0
Models
SEIR, eq. 20 101.7 173.0 258.1 17.1 49.6 105.4 17.9 19.8 30.8 0.0
Mean of Relative Errors in %
Software IQM SciML AMIGO2 Parameter
Estimation.jl
Search Range [0,1] [0,2] [0,3] [0,1] [0,2] [0,3] [0,1] [0,2] [0,3] Any
Harmonic, eq. 2 66.3 68.7 101.9 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Van der Pol, eq. 4 7.5 0.0 9.6 0.0 0.0 0.0 0.0 0.0 0.0 0.0
FitzHugh-Nagumo, eq. 6 82.4 131.1 210.3 8.3 15.9 20.4 0.0 0.0 0.2 0.0
HIV, eq. 8 85.6 104.2 144.7 14.8 36.9 102.9 18.5 43.2 28.4 0.0
Mammillary 3, eq. 10 76.1 104.3 128.2 11.9 21.5 24.7 0.0 0.0 0.0 0.0
Lotka-Volterra, eq. 12 72.2 75.4 75.4 54.7 76.9 111.7 13.6 112.6 29.7 0.0
Crauste, eq. 14 99.6 125.2 179.7 47.3 72.9 226.4 2.6 105.7 51.7 0.0
12
Biohydrogenation, eq. 16 93.7 184.0 130.3 77.7 151.1 306.6 17.8 13.5 28.3 0.0
Mammillary 4, eq. 18 94.1 109.6 105.3 66.2 67.7 118.8 29.8 56.7 59.0 0.1
Models
SEIR, eq. 20 132.3 230.8 338.9 18.4 60.6 118.1 25.9 40.7 50.8 0.0
Table 1: The above two tables show the median and the mean of the Root Mean Square Relative Errors (RMSRE) over 10 randomly generated data sets. For the details on how the data sets are generated and how the errors are computed, see Subsections 5.25.3 and 5.4. The last three models are grouped separately from the rest to indicate that the model has parameters that are only locally identifiable.
In this section, we study the performance (in particular robustness) of several software packages on a collection of benchmarks models listed in section 6.
5.1 Software Packages
We ran the following software packages:
  • IQM [56]
    It uses a single shooting approach with a local derivative-free optimizer (Nelder-Mead, as specified by symplexIQM).
  • SciML [50, 51]
    It is intentionally a naive single shooting implementation which serves as a baseline for one of the most common parameter estimation approaches. While such a package can easily be improved using methods like multiple shooting and global optimizers, this setup demonstrates the expected results from the standard simple approach. A polyalgorithm mixing ADAM with BFGS is used to demonstrate the results of a common gradient-based optimization scheme.
  • AMIGO2 [2]
    It uses a multi-start global optimization approach. This approach can always be tuned to achieve correct results by choosing sufficient bounds/tolerances and population points, though not all of these values are typically known in advance.
  • ParameterEstimation.jl (this work)
We could not run the following software packages: Data2Dynamics [54], COPASI [29], and IQR [^1] https://iqrtools.intiquan.com//IQDesktop [^2] https://iqdesktop.intiquan.com/book/ because of technical difficulties in installation.
5.2 Settings used for Software Packages
IQM, SciML and AMIGO2 require search ranges or initial guesses for the parameters and initial conditions or certain additional settings to be specified by the user that pertain specifically to the optimization task performed by each program, for instance, maximal # of iterations for gradient-based optimization, specific loss/error functions to be optimized.
ParameterEstimation.jl is built on top of existing algorithms for solving polynomial systems and interpolating data points. Their implementations involve some settings.
For studying performance of the software systems, we used the following settings:
  • For IQM, we set the maximal number of function evaluations at 200,000 and used a simplexIQM solver.
  • For SciML, we set the number of function evaluation at 200,000 and set the learning rate at 0.01.
  • For AMIGO2, we set the maximal number of function evaluations at 200,000, set the IVP solver integration tolerances at $10^{-13}$ and set the solver tolerances to $10^{-13}$ and used a nonlinear least squares solver, which meets or exceeds the settings suggested by Julio Banga [^3] Private communication, May-June 2023. Due to the non-deterministic nature of choosing starting points in the global optimization used by AMIGO2, multiple runs with the same example and settings can produce different estimates (possibly with larger or smaller error). We only ran it once and reported the results.
  • For ParameterEstimation.jl, default options were used for most settings. The default polynomial solver is HomotopyContinuation.jl, with a tolerance of $10^{-12}$. The default setting is to use a fixed set of interpolation schemes including AAA, Floater-Hormann of orders (3,6,8), and Fourier interpolation (though, for all of the models in this paper, AAA interpolation alone is sufficient.)
5.3 Test Data Generation
For all models and all software packages, the test data were generated uniformly as follows. We chose 21 time points equally spaced over $[-0.5,0.5]$ and also chose the reference values of the parameters and initial conditions randomly from the interval $[0.1,0.9]$ using the random.uniform function in Python’s numpy package. For reproducibility, we used random.seed(0) at the start of choosing the 10 data sets for each system. In case of errors (e.g., due to singularity), the set was discarded and the next set of random values were generated instead. These uniform random choices were made for the sake of studying/comparing the robustness of the approaches in isolation, independent of the physical systems underlying the models. The random values were at least 0.1 chosen away from 0.0 and 1.0 to avoid potential degeneracy.
For each model and for each software package, the input script (containing the model description, the generated data and settings) was generated automatically.
5.4 Estimation Error
To evaluate the quality of the estimation, for each data set, we computed the Root Mean Square Relative Error defined as
$RMSRE=\sqrt{\frac{1}{\#P}\sum_{\begin{subarray}{c}p\end{subarray}\in P}\left(\frac{p_{true}-p_{estimated}}{p_{true}}\right)^{2}}\times 100,$
where $P$ denotes the identifiable444 For the Biohydrogenation model, we excluded the errors from the estimates for the initial value of $x_{7}$, which is not identifiable. set of the system parameters and the initial values of the state variables in the input system. The appendix provides RMSRE for each model, search interval, and random data set. Table 1 summarizes them using median and mean.
5.5 Discussion
We make the following observations from the test results.
  • Relative error: For most of the scenarios considered here, our system was able to outperform in estimation error. In all but one instance, we observed that our method outperformed the others with regard to estimation error.
  • Multiple solutions: Our program’s advantage is its ability to recover multiple solutions if appropriate (local identifiability) while optimization-based programs may miss a valid value of a parameter.
One of the most notable improvements of our approach over the shooting-based methods is the lack of required hyperparameters or the necessity of choosing the search range for the parameter values. In contrast, e.g., in AMIGO2, one must choose lower and upper bounds for the ODE parameters and initial conditions. In particular, in the Crauste Model, eq. 14, if the upper bound of the range for all parameters is increased from 1.0 to 2.0, the median (mean) of RMSRE’s increased from 0.0% (2.6%) to 77.6% (105.7%). However, if the range is further increased to 3.0, the median (mean)of RMSRE’s decreased to 24.5% (51.7%). (See Random Data 3, 5, and 8-10 in the Crauste table of Appendix for specific instances of this behavior.) This shows the non-monotonic behavior of tuning the approaches that rely on global optimization, which can add difficulty for scientists to accurately use the methods in practice. This is not a criticism of these approaches but is rather a note that the potential users have to be aware of.
Our parameter estimation approach does not require any hyperparameters while giving all solutions (e.g., if some parameters are only locally but not globally identifiable), making it easier for a practicing scientist to quickly understand the full landscape of solutions.
6 List of Differential Models
In this section, we present the ODE systems used in our benchmarks.
Harmonic Oscillator Model
It is a toy model for harmonic oscillators without damping.
$\begin{cases}\dot{x_{1}}=-a\,x_{2}\\ \dot{x_{2}}=\frac{1}{b}\,x_{1}\end{cases}$ (2)
$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\end{cases}$ (3)
Van der Pol Oscillator Model
It models harmonic oscillators with nonlinear damping [61], which was introduced to study oscillations in electronic circuits, but was found to be useful for various other phenomena such as vocal fold [40].
$\begin{cases}\dot{x_{1}}=ax_{2}\\ \dot{x_{2}}=-x_{1}-b(x_{1}^{2}-1)x_{2}\end{cases}$ (4)
$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\end{cases}$ (5)
FitzHugh-Nagumo Model
It models (two-dimensional simplification of) spike generation in squid giant axons [22, 44].
$\begin{cases}\dot{V}=g\,\left(V-\frac{V^{3}}{3}+R\right)\\ \dot{R}=\frac{1}{g}\,\left(V-a+b\,R\right)\end{cases}$ (6)
$\begin{cases}y_{1}=V\end{cases}$ (7)
HIV Dynamics Model
It models HIV infection dynamics during interaction with the immune system over the course of various treatments. It is based on  [66, equations (6)] and the paragraph that follows.
$\begin{cases}\dot{x}=\lambda-d\,x-\beta\,x\,v\\ \dot{y}=\beta\,x\,v-a\,y\\ \dot{v}=k\,y-u\,v\\ \dot{w}=c\,x\,y\,w-c\,q\,y\,w-b\,w\\ \dot{z}=c\,q\,y\,w-h\,z\end{cases}$ (8)
$\begin{cases}y_{1}=w,~{}y_{2}=z\\ y_{3}=x,~{}y_{4}=y+v\end{cases}$ (9)
Mammillary 3-compartment Model
It models pharmacokinetic processes, such as protein metabolism in organism [8, 30].
$\begin{cases}\dot{x_{1}}=-(a_{21}+a_{31}+a_{01})\,x_{1}+a_{12}\,x_{2}+a_{13}\,x_{3}\\ \dot{x_{2}}=a_{21}\,x_{1}-a_{12}\,x_{2}\\ \dot{x_{3}}=a_{31}\,x_{1}-a_{13}\,x_{3}\end{cases}$ (10)
$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\end{cases}$ (11)
Lotka-Volterra Model
It models interactions between two species (a predator and a prey) in an ecosystem  [39, 64].
$\begin{cases}\dot{r}=k_{1}\,r-k_{2}\,r\,w,\\ \dot{w}=k_{2}\,r\,w-k_{3}\,w\end{cases}$ (12)
$\begin{cases}y_{1}=r\end{cases}$ (13)
Crauste Model
It models the behavior of CD8 T-cells. It is based on the Crauste system [26, Section 4] originally described in [17].
$\begin{cases}\dot{N}=-\mu_{N}\,N-\delta_{NE}\,N\,P\\ \dot{E}=\delta_{NE}\,N\,P-\mu_{EE}\,E^{2}-\delta_{EL}\,E+\rho_{E}\,E\,P\\ \dot{S}=\delta_{EL}\,S-S\,\delta_{LM}-\mu_{LL}\,S^{2}-\mu_{LE}\,E\,S\\ \dot{M}=\delta_{LM}\,S-\mu_{M}\,M\\ \dot{P}=\rho_{P}\,P^{2}-\mu_{P}\,P-\mu_{PE}\,E\,P-\mu_{PL}\,S\,P\end{cases}$ (14)
$\begin{cases}y_{1}=N,~{}y_{2}=E\\ y_{3}=S+M,y_{4}=P\end{cases}$ (15)
Biohydrogenation Model
It models the kinetic processes involved in biohydrogenation of non-esterified fatty acids  [43, 42].
$\begin{cases}\dot{x_{4}}=-\frac{k5\,x_{4}}{k_{6}+x_{4}}\\ \dot{x_{5}}=\frac{k_{5}\,x_{4}}{k_{6}+x_{4}}-\frac{k_{7}\,x_{5}}{k_{8}+x_{5}+x_{6}}\\ \dot{x_{6}}=\frac{k_{7}\,x_{5}}{k_{8}+x_{5}+x_{6}}-\frac{k_{9}\,x_{6}\,(k_{10}-x_{6})}{k_{10}}\\ \dot{x_{7}}=\frac{k_{9}\,x_{6}\,(k_{10}-x_{6})}{k_{10}}\end{cases}$ (16)
$\begin{cases}y_{1}=x_{4}\\ y_{2}=x_{5}\end{cases}$ (17)
Mammillary 4-Compartment Model
It comes from examples used in the analysis of DAISY [6] identifiability software. It represents a 4-compartment system.
$\begin{cases}\dot{x_{1}}=-k_{01}\,x_{1}+k_{12}\,x_{2}+k_{13}\,x_{3}+\\ \hskip 35.00005ptk_{14}x_{4}-k_{21}\,x_{1}-k_{31}\,x_{1}-k_{41}\,x_{1}\\ \dot{x_{2}}=-k_{12}\,x_{2}+k_{21}\,x_{1}\\ \dot{x_{3}}=-k_{13}\,x_{3}+k_{31}\,x_{1}\\ \dot{x_{4}}=-k_{14}\,x_{4}+k_{41}\,x_{1}\\ \end{cases}$ (18)
$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\\ y_{3}=x_{3}+x_{4}\end{cases}$ (19)
SEIR Model
It models an epidemic that involves compartments related to disease progression stages with prevalence observations [59].
$\begin{cases}\dot{S}=-b\,S\,I/N\\ \dot{E}=b\,S\,I/N-\nu\,E\\ \dot{I}=\nu\,E-a\,I\\ \dot{N}=0\end{cases}$ (20)
$\begin{cases}y_{1}=I\\ y_{2}=N\end{cases}$ (21)
Acknowledgements
The authors are grateful to the CCiS at Queens College and CIMS NYU for the computational resources, to Julio Banga and Eva Balsa-Canto for extensive help and feedback on the use of AMIGO2. The authors are also grateful to Xinglong Zhou and Jianing Qi for initial explorations of the role of polynomial interpolation in this problem. The authors are also grateful to Henning Schmidt, developer of IQRTools/IQDesktop for numerous discussions while we tried to install the software. We thank Gleb Pogudin for helpful comments on our algorithm and additional information on other parameter estimation methods.
This work was supported by the National Science Foundation [DMS-1760448, CCF-1708884, CCF-2212460, CCF-2212461, CCF-2212462, CCF-1813340], City University of New York [PSC-CUNY #65605-00 53, This work was supported by the National Science Foundation [DMS-1760448, CCF-1708884, CCF-2212460, CCF-2212461, CCF-2212462] and City University of New York [PSC-CUNY #65605-00 53, #66551-00 54], and DARPA [HR00112290091]. This research was, in part, funded by the U.S. Government. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. Approved for Public Release, Distribution Unlimited.

### 17 Dec 2023 v3

@@
 We present a new approach for estimating parameters in rational ODE models from given (measured) time series data.
-In a typical existing approach, one first tries to make a good initial guess for the parameter values.
+In typical existing approaches, an initial guess for the parameter values is made from a given search interval.
 Then, in a loop, the corresponding outputs are computed by solving the ODE numerically, followed by computing the error from the given time series data.
@@
 Otherwise, heuristics/theories are used to possibly improve the guess and continue the loop.
-A downside of this approach is non-robustness, as there are no guarantees for the result of the loop iterations to be predictably close to the true parameter values.
+These approaches tend to be non-robust in the sense that their accuracy depend on the search interval and the true parameter values; furthermore, they cannot handle the case where the parameters are locally identifiable.
 In this paper, we propose a new approach, which does not suffer from the above non-robustness.
-In particular, it does not require making good initial guesses for the parameter values.
-Instead, it uses differential algebra, interpolation of the data using rational functions, and multivariate polynomial system solving, and has a potential for a complete user control over the error of the estimation (the actual error analysis is left for the future research).
+In particular, it does not require making good initial guesses for the parameter values or specifying search intervals.
+Instead, it uses differential algebra, interpolation of the data using rational functions, and multivariate polynomial system solving.
 We also compare the performance of the resulting software with several other estimation software packages.
@@
 1.2 State of the art
-Due to its importance, this problem has been the subject of intensive research efforts that have yielded various theories and methods.
+Due to its importance, this problem has been the subject of intensive research efforts that have yielded various theories and approaches.
 They can be roughly categorized into three approaches: shooting, two-stage, and algebraic.
@@
 This popular approach has seen many software implementations; to list a few: AMIGO2, COPASI, Data2Dynamics, SBtoolbox2/IQM; see [3, 5, 38, 65, 58, 56, 36, 29, 23, 21, 2, 54, 25, 63, 19].
-Many studies have further developed the shooting methods to improve robustness, adding techniques such as multiple shooting [10, 1, 60], automatic differentiation and adjoint techniques for more accurate derivatives [41, 46, 33], and landscape smoothing techniques like the prediction-error method [34], all of which help the optimization techniques avoid local minima.
+Many studies have further developed the shooting approaches to improve robustness, adding techniques such as multiple shooting [10, 1, 60], automatic differentiation and adjoint techniques for more accurate derivatives [41, 46, 33], and landscape smoothing techniques like the prediction-error method [34], all of which help the optimization techniques avoid local minima.
 Two-stage approaches (also known as collocation, smoothing, or principle differential analysis) avoid numerically solving the ODE by fitting collocation polynomials to the data to obtain predictions of the data which are then directly fit against the data [52, 14].
-As the result, these methods are generally much more efficient.
-Previous studies have used various mathematical objects such as B-splines [13, 48, 35] and local least-squares [37, 55, 53, 15].
-One downside to two-stage methods is requiring observation of all states.
+As the result, these approaches are generally much more efficient.
+Previous studies have used various objects such as B-splines [13, 48, 35] and local least-squares [37, 55, 53, 15].
+One downside to two-stage approaches is requiring observation of all states.
 Furthermore, it has been noted that the error introduced through the derivative estimation tends to lead to less accurate final results in real-world scenarios with less dense data [48, 14].
-These studies suggest finishing a parameter estimation by using the output of a two-stage method as the initial condition to a simulation-based method.
+These studies suggest finishing a parameter estimation by using the output of a two-stage approach as the initial condition to a simulation-based approach.
 Algebraic approaches tackle rational differential models by exploiting theoretical results from differential algebra.
@@
 Finally, the paper  [32] investigates the parameter manifold in the case where the differential model is structurally non-identifiable.
-1.3 Our approach and novelty
-Our approach could roughly be classified as a two stage approach.
-However, it is substantially different from the previous two stage approaches, in that it computes optimal parameters robustly.
-For instance, it can reliably handle even degenerate cases such as when the parameters are only locally identifiable.
+1.3 The proposed approach and novelty
+The proposed approach could roughly be classified as a two stage approach.
+However, it is substantially different from the previous two stage approaches, in that it estimates parameters robustly, in the following sense:
+  1. 1.
+    The proposed approach produces good estimates no matter what the specified search intervals are. In fact, it does not even require specifying any search interval.
+  2. 2.
+    The proposed approach produces good estimates no matter where the true parameter values are.
+  3. 3.
+    The proposed approach produces good estimates even when the parameters are not globally identifiable, as long as they are locally identifiable.
 It achieves the robustness by judiciously combining:
@@
   • overdetermined polynomial system solving by squaring the system and filtering the solution set.
-Our approach has the following novel features:
-  1. 1.
-    Our approach does not involve making initial guesses or global optimization; it instead relies on differential algebraic properties of the ODE model to propose candidates for parameter values.
-  2. 2.
-    Out of the possible solution candidates, the procedure computes the best fit for the given data including cases in which such fit may not be unique.
-That is, accounting for possible local identifiability of the model, we report all valid parameter values.
-The user can then quickly filter out some (or all) of the not possible parameter values if the users has additional information about these values such as ranges they take values in.
-As such, this technique can be used in isolation for dense data, or together with shooting methods by providing more robust initial conditions for sparse data.
+As such, this technique can be used in isolation for dense data, or together with shooting approaches by providing more robust initial conditions for sparse data.
 1.4 Software Implementation
-We implemented our approach into a Julia package called “ParameterEstimation.jl”. All code and benchmarking data are available at https://github.com/iliailmer/ParameterEstimation.jl.
+We implemented the proposed approach into a Julia package called “ParameterEstimation.jl”. All code and benchmarking data are available at https://github.com/iliailmer/ParameterEstimation.jl.
 2 Problem Statement
-In this section, we precisely state the problem solved by our approach and illustrate it using a toy example. We also show how to enter/solve the problem using our soft implementation.
+In this section, we precisely state the problem solved by the proposed approach and illustrate it using a toy example. We also show how to enter/solve the problem using our soft implementation.
 Input:
@@
 Initial Condition(s): x(t) = 1.000
-3 Our Approach
-We will briefly describe our approach by first giving a flow chart of the steps.
+3 The proposed Approach
+We will briefly describe the proposed approach by first giving a flow chart of the steps.
 Input: $\begin{array}[]{ll}\text{Model:}&\left\{\begin{array}[]{ll}x^{\prime}&=-\mu x\\ y&=x^{2}+x\end{array}\right.\\ \text{Data:}&\begin{array}[]{l}\left(0.000,2.000\right),\left(0.333,1.563\right)\\ \left(0.666,1.229\right),\left(1.000,0.974\right)\end{array}\end{array}$ of values for the parameters = 1 Differentiated system: $y$ $=x^{2}+x$ $x^{\prime}$ $=-\mu x$ $y^{\prime}$ $=2xx^{\prime}+x^{\prime}$ $x^{\prime\prime}$ $=-\mu x^{\prime}$ $y^{\prime\prime}$ $=2\left(x^{\prime}x^{\prime}+xx^{\prime\prime}\right)+x^{\prime\prime}$ Polynomial system: 2.002.00 2.002.00 $\approx x_{0}^{2}+x_{0}$ $x_{1}$ $=-\mu x_{0}$ $-1.50$ $\approx 2x_{1}x_{0}+x_{1}$ $x_{2}$ $=-\mu x_{1}$ 1.221.22 1.221.22 $\approx 2\left(x_{1}^{2}+x_{0}x_{2}\right)+x_{2}$ Solutions: $S_{1}=\left\{\begin{array}[]{ll}\mu&=0.499\\ x_{0}&=1.000\end{array}\right.S_{2}=\left\{\begin{array}[]{ll}\mu&=0.249\\ x_{0}&=-2.000\end{array}\right.$ Errors: $\begin{array}[]{ll}S_{1}\Rightarrow{\bf 6.87\cdot 10^{-4}}&S_{2}\Rightarrow 2.22\cdot 10^{-2}\end{array}$ Output: $\mu=0.499,~{}x(0)=1.000$ 1) Find # values for parameters 2) Differentiate 3) Use the data 4) Find all solutions5) Compute errors6) Select
@@
     Use the model $\Sigma$ to find the number of values for parameters (see [27, 28] for details, cf. [4]).
-    Toy example: There is only one value for the parameters.
+    Toy example:
+    There is only one value for the parameters.
   2. 2.
@@
 Experimental data show that AAA rational interpolation interpolates smooth functions in such a way that, for a wide variety of data outputs from differential equations, even higher order derivatives are well interpolated.
-In our program, we use AAA as a default interpolation scheme, but allow iterating over multiple schemes in a parallel fashion to select the best one.
+In our software, we use AAA as a default interpolation scheme, but allow iterating over multiple schemes in a parallel fashion to select the best one.
 Once interpolation is done, we apply high-order automatic differentiation provided by the TaylorSeries.jl [7] and ForwardDiff.jl packages to the interpolated function to estimate derivative values of higher order.
@@
 Current Limitations
-We list a few limitations of the current implementation:
-  • As a two-stage method, the issues common to this class of methods apply.
+We list a few limitations of the current preliminary implementation:
+  • As a two-stage approach, the issues common to this class of approaches apply.
 The data must be time series data which is sufficiently dense for the rational polynomial fitting to give accurate derivative approximations.
-  • Our program supports rational ODEs models (the right-hand sides are rational functions), while software such as AMIGO2 and IQM also support ODE models that are not necessarily rational.
+  • Our software supports rational ODEs models (the right-hand sides are rational functions), while software such as AMIGO2 and IQM also support ODE models that are not necessarily rational.
   • We use SIAN that is based on a Monte-Carlo algorithm that returns the bound on number of solutions with high probability, though detailed analysis is still ongoing.
@@
 5.2 Settings used for Software Packages
-IQM, SciML and AMIGO2 require search ranges or initial guesses for the parameters and initial conditions or certain additional settings to be specified by the user that pertain specifically to the optimization task performed by each program, for instance, maximal # of iterations for gradient-based optimization, specific loss/error functions to be optimized.
+IQM, SciML and AMIGO2 require search ranges or initial guesses for the parameters and initial conditions or certain additional settings to be specified by the user that pertain specifically to the optimization task performed by each software package, for instance, maximal # of iterations for gradient-based optimization, specific loss/error functions to be optimized.
 ParameterEstimation.jl is built on top of existing algorithms for solving polynomial systems and interpolating data points. Their implementations involve some settings.
@@
 5.5 Discussion
-We make the following observations from the test results.
-  • Relative error: For most of the scenarios considered here, our system was able to outperform in estimation error.
-In all but one instance, we observed that our method outperformed the others with regard to estimation error.
-  • Multiple solutions: Our program’s advantage is its ability to recover multiple solutions if appropriate (local identifiability) while optimization-based programs may miss a valid value of a parameter.
-One of the most notable improvements of our approach over the shooting-based methods is the lack of required hyperparameters or the necessity of choosing the search range for the parameter values.
-In contrast, e.g., in AMIGO2, one must choose lower and upper bounds for the ODE parameters and initial conditions.
-In particular, in the Crauste Model, eq.
-14, if the upper bound of the range for all parameters is increased from 1.0 to 2.0, the median (mean) of RMSRE’s increased from 0.0% (2.6%) to 77.6% (105.7%).
-However, if the range is further increased to 3.0, the median (mean)of RMSRE’s decreased to 24.5% (51.7%).
-(See Random Data 3, 5, and 8-10 in the Crauste table of Appendix for specific instances of this behavior.) This shows the non-monotonic behavior of tuning the approaches that rely on global optimization, which can add difficulty for scientists to accurately use the methods in practice.
-This is not a criticism of these approaches but is rather a note that the potential users have to be aware of.
-Our parameter estimation approach does not require any hyperparameters while giving all solutions (e.g., if some parameters are only locally but not globally identifiable), making it easier for a practicing scientist to quickly understand the full landscape of solutions.
+The inspection of the test results in Table 1 and Appendix shows that the proposed approach outperforms the others with regard to estimation error in all but one instance.
+Furthermore, the proposed approach is robust in the following aspects:
+Search Interval: See Table 1.
+Note that IQM, SciML and AMIGO2 (based on shooting approaches) require the user to specify a search interval.
+Note also that their relative errors are dependent on search interval, sometimes even non-monotonically.
+For an example, see the relative error of AIMOG2 on the Crauste Model.
+If the upper bound of the search interval is increased from 1.0 to 2.0, the median (mean) of the relative error is increased from 0.0% (2.6%) to 77.6% (105.7%).
+However, if the upper bound is increased to 3.0, the median (mean) of the relative error is decreased to 24.5% (51.7%).
+See Random Data 3, 5, 9 and 10 in the Crauste table of Appendix for specific instances of this behavior.
+This non-monotonic behavior could add difficulty to the users in specifying a suitable search interval.
+The proposed approach’s relative error is, by design, independent of search interval, because it does not require/use search interval.
+Parameter values: See the tables in Appendix.
+Note that the relative errors of IQM, SciML and AMIGO2 are dependent on the values of the parameters (randomly sampled from [0.1, 0.9]), sometime drastically.
+For an example, see the Crauste table.
+AMIGO2 has essentially no error for some random values of parameters but huge error for the others.
+The proposed approach’s relative error is practically independent of the values of the parameters.
+Multiple solutions: See Table 1.
+The first seven models are globally identifiable (hence with unique solution for the parameters).
+The last three models contain parameters that are only locally identifiable (hence with multiple solutions for the parameter values).
+Note that the relative errors of IQM, SciML and AMIGO2 are significantly larger for locally identifiable models.
+It is because the optimization-based approaches may miss an intended value of a parameter.
+The proposed approach’s relative error is practically independent of identifiability, as long as it is at least locally identifiable.
+In fact, it finds all the solutions, making it easier for the users to quickly understand the full landscape of solutions.
 6 List of Differential Models
-In this section, we present the ODE systems used in our benchmarks.
+In this section, we present the ODE systems used in the benchmarks.
 Harmonic Oscillator Model
@@
 $\begin{cases}\dot{x_{1}}=-a\,x_{2}\\ \dot{x_{2}}=\frac{1}{b}\,x_{1}\end{cases}$ (2)
-$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\end{cases}$ (3)
+$\begin{cases}y_{1}=x_{1},\\ y_{2}=x_{2}\end{cases}$ (3)
 Van der Pol Oscillator Model
@@
 $\begin{cases}\dot{x_{1}}=ax_{2}\\ \dot{x_{2}}=-x_{1}-b(x_{1}^{2}-1)x_{2}\end{cases}$ (4)
-$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\end{cases}$ (5)
+$\begin{cases}y_{1}=x_{1},\\ y_{2}=x_{2}\end{cases}$ (5)
 FitzHugh-Nagumo Model
@@
 $\begin{cases}\dot{x}=\lambda-d\,x-\beta\,x\,v\\ \dot{y}=\beta\,x\,v-a\,y\\ \dot{v}=k\,y-u\,v\\ \dot{w}=c\,x\,y\,w-c\,q\,y\,w-b\,w\\ \dot{z}=c\,q\,y\,w-h\,z\end{cases}$ (8)
-$\begin{cases}y_{1}=w,~{}y_{2}=z\\ y_{3}=x,~{}y_{4}=y+v\end{cases}$ (9)
+$\begin{cases}y_{1}=w,\\ y_{2}=z,\\ y_{3}=x,\\ y_{4}=y+v\end{cases}$ (9)
 Mammillary 3-compartment Model
@@
 $\begin{cases}\dot{x_{1}}=-(a_{21}+a_{31}+a_{01})\,x_{1}+a_{12}\,x_{2}+a_{13}\,x_{3}\\ \dot{x_{2}}=a_{21}\,x_{1}-a_{12}\,x_{2}\\ \dot{x_{3}}=a_{31}\,x_{1}-a_{13}\,x_{3}\end{cases}$ (10)
-$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\end{cases}$ (11)
+$\begin{cases}y_{1}=x_{1},\\ y_{2}=x_{2}\end{cases}$ (11)
 Lotka-Volterra Model
@@
 $\begin{cases}\dot{N}=-\mu_{N}\,N-\delta_{NE}\,N\,P\\ \dot{E}=\delta_{NE}\,N\,P-\mu_{EE}\,E^{2}-\delta_{EL}\,E+\rho_{E}\,E\,P\\ \dot{S}=\delta_{EL}\,S-S\,\delta_{LM}-\mu_{LL}\,S^{2}-\mu_{LE}\,E\,S\\ \dot{M}=\delta_{LM}\,S-\mu_{M}\,M\\ \dot{P}=\rho_{P}\,P^{2}-\mu_{P}\,P-\mu_{PE}\,E\,P-\mu_{PL}\,S\,P\end{cases}$ (14)
-$\begin{cases}y_{1}=N,~{}y_{2}=E\\ y_{3}=S+M,y_{4}=P\end{cases}$ (15)
+$\begin{cases}y_{1}=N,\\ y_{2}=E,\\ y_{3}=S+M,\\ y_{4}=P\end{cases}$ (15)
 Biohydrogenation Model
@@
 $\begin{cases}\dot{x_{4}}=-\frac{k5\,x_{4}}{k_{6}+x_{4}}\\ \dot{x_{5}}=\frac{k_{5}\,x_{4}}{k_{6}+x_{4}}-\frac{k_{7}\,x_{5}}{k_{8}+x_{5}+x_{6}}\\ \dot{x_{6}}=\frac{k_{7}\,x_{5}}{k_{8}+x_{5}+x_{6}}-\frac{k_{9}\,x_{6}\,(k_{10}-x_{6})}{k_{10}}\\ \dot{x_{7}}=\frac{k_{9}\,x_{6}\,(k_{10}-x_{6})}{k_{10}}\end{cases}$ (16)
-$\begin{cases}y_{1}=x_{4}\\ y_{2}=x_{5}\end{cases}$ (17)
+$\begin{cases}y_{1}=x_{4},\\ y_{2}=x_{5}\end{cases}$ (17)
 Mammillary 4-Compartment Model
@@
 $\begin{cases}\dot{x_{1}}=-k_{01}\,x_{1}+k_{12}\,x_{2}+k_{13}\,x_{3}+\\ \hskip 35.00005ptk_{14}x_{4}-k_{21}\,x_{1}-k_{31}\,x_{1}-k_{41}\,x_{1}\\ \dot{x_{2}}=-k_{12}\,x_{2}+k_{21}\,x_{1}\\ \dot{x_{3}}=-k_{13}\,x_{3}+k_{31}\,x_{1}\\ \dot{x_{4}}=-k_{14}\,x_{4}+k_{41}\,x_{1}\\ \end{cases}$ (18)
-$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\\ y_{3}=x_{3}+x_{4}\end{cases}$ (19)
+$\begin{cases}y_{1}=x_{1},\\ y_{2}=x_{2},\\ y_{3}=x_{3}+x_{4}\end{cases}$ (19)
 SEIR Model
@@
 $\begin{cases}\dot{S}=-b\,S\,I/N\\ \dot{E}=b\,S\,I/N-\nu\,E\\ \dot{I}=\nu\,E-a\,I\\ \dot{N}=0\end{cases}$ (20)
-$\begin{cases}y_{1}=I\\ y_{2}=N\end{cases}$ (21)
+$\begin{cases}y_{1}=I,\\ y_{2}=N\end{cases}$ (21)
 Acknowledgements
-The authors are grateful to the CCiS at Queens College and CIMS NYU for the computational resources, to Julio Banga and Eva Balsa-Canto for extensive help and feedback on the use of AMIGO2.
-The authors are also grateful to Xinglong Zhou and Jianing Qi for initial explorations of the role of polynomial interpolation in this problem.
-The authors are also grateful to Henning Schmidt, developer of IQRTools/IQDesktop for numerous discussions while we tried to install the software.
-We thank Gleb Pogudin for helpful comments on our algorithm and additional information on other parameter estimation methods.
+We, the authors, are grateful to Julio Banga and Eva Balsa-Canto for extensive feedback on the use of AMIGO2, to Henning Schmidt, developer of IQRTools/IQDesktop for numerous discussions on installation of the software, to Gleb Pogudin for helpful feedback on the proposed algorithm and information on other parameter estimation approaches, to Xinglong Zhou and Jianing Qi for initial explorations of the role of polynomial interpolation, and finally to the CCiS at Queens College and CIMS NYU for the computational resources.
 This work was supported by the National Science Foundation [DMS-1760448, CCF-1708884, CCF-2212460, CCF-2212461, CCF-2212462, CCF-1813340], City University of New York [PSC-CUNY #65605-00 53, This work was supported by the National Science Foundation [DMS-1760448, CCF-1708884, CCF-2212460, CCF-2212461, CCF-2212462] and City University of New York [PSC-CUNY #65605-00 53, #66551-00 54], and DARPA [HR00112290091].